OpenStudy (anonymous):

The population of a town was 34,000 in 1940 and 62,000 in 2010. Assume the rate of population increase is constant. What will the population be in 2015?

OpenStudy (anonymous):

from 34,000 to 62,000 is an increase of 28,000. 28,000 divided by the years, is 2010 - 1940 = 70. so then you get, 28,000 / 70 = 400 so, per year there was an increase of 400 people. So then from 2010 to 2015 would be 2015-2015 = 5 years so, 62,000 + ( 5 * 400) 62,000 + 2000 = 64,000

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

youre welcome :)